For the next balance equation:

ki ES E S k.11

a) Calculate the concentration of [S] at a time in the reaction where [ES] = 0.09 mM if [S]0 was 2.55 mM.

b) What was [E]0 if at the same time [E] was 0.02 mM?

c) If the reaction is in equilibrium, calculate the equilibrium constant value for the reaction and the ratio between k1 and k-1.
Answer #1

A) 2.55 - 0.09 =2.460

B) 0.02+0.09= 0.11

C) k= 0.09/2.46×0.02

=0.036

K={ES}/{E}×{S}......1

K-1={E}{S}/{ES}......2

K/k-1   =k2=( 0.036)2

= 0.0013